Mary is 12 years and her father is 42 years old
how old will her father be we he is 4 times as old as mary

Found 2 solutions by macston, stanbon:
Answer by macston(5194)

M=Mary=F-30; F=Father=M+30;
F=4M Substitute for M
F=4(F-30)
F=4F-120 Subtract F from each side.
0=3F-120 Add 120 to each side.
120=3F Divide each side by 3
40=F
ANSWER Mary's father was 40 when he was 4 times as old as Mary.

Answer by stanbon(75887)

Mary is 12 years and her father is 42 years old
how old was her father when he was 4 times as old as mary?
----
Equations:
f = m + 30
42+x = 4(12+x)
------
42+x = 48+4x
3x = -6
x = -2
----
When father was 40, mary was 10.
----------------------------------
Cheers,
Stan H.
-------------